How do you put factor form into vertex form?
example: y=(x^2+2)(x^2+5)

Answer by RAY100(1637)

y = ( x^2 +2) ( X^2 +5 )
.
y = x^4 +7x^2 +10
.
y-10 = x^4 + 7 x^2 +12.25 -12.25,,,,,,,,,remember just use (7/2)^2 = 12.25
.
y-10 +12.25 = x^4 +7x^2 +12.25
.
( y +2.25 ) = ( x^2 + 3.5 )^2
.
checking
x=0, both original and answer, y=10,,,,ok